# Lie Neurons: Adjoint-Equivariant Neural Networks for Semisimple Lie Algebras

### v2

@@ 5.1 Experiments on $\mathfrak{so}(3)$
As discussed in Section 4.6, our network when applied on $\mathfrak{so}(3)$ resembles Vector Neurons (Deng et al., 2021), but contains more flexible component layers. We conduct experiments on BCH formula regression and dynamics learning, which highlight the benefit of being able to process 3-D vectors as $\mathfrak{so}(3)$ elements. We also conduct experiments on point cloud registration, where Vector Neurons have been shown to work (Zhu et al., 2022). We show that our method has comparable performance in these tasks and even outperform in some cases.
@@ 5.1.1 Baker–Campbell–Hausdorff Formula
The Baker-Cambell-Hausdorff (BCH) formula provides a way to compute the product of two exponentials of elements in a Lie algebra locally (Hall, 2013). For $X,Y,Z\in\mathfrak{g}$, and
@@
This formula is widely used in robotics and control (Yoon et al., 2023; Chauchat et al., 2018; Kobilarov & Marsden, 2011). However, the higher-order terms are often discarded in most applications, which can result in a significant drop in accuracy. An accurate modeling of the BCH formula can have great benefits in the above field. In light of this, we use the Lie Neurons to model the BCH formula in $\mathfrak{so}(3)$ in this experiment. We note that the BCH formula is adjoint equivariant. To assess the performance of the proposed method, we compare the Lie Neurons with the Vector Neurons and an MLP.
In order to train the network, we generate $10,000$ data points for both $X$ and $Y$ and use (19) to generate the ground truth value. We limit the angle to $[0,\pi)$ to ensure an injective exponential function. We design the loss function to be:
@@
where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm, and $I$ is the identity matrix. The network architecture used in this experiment can be found in Figure 5 in the Appendix. In addition to the Frobenius norm, we also report the log error, which is defined as:
$E_{log}=||\log(e^{X}e^{Y}e^{-f(X,Y)})||,$ (22)
where |||||$\cdot$||| | ⋅ | | is the standard vector norm on $\mathbb{R}^{3}$.
The results are presented in Table 1. Our method clearly outperforms the baselines. While the MLP is able to produce acceptable results on the non-conjugated test set, its performance suffers from a significant drop when the inputs are adjoint-transformed. The Vector Neurons is unable to converge in this experiment. We observe a similar result when we only use the ReLU layer as the activation function for the Lie Neurons. We conjecture this to be because of the inherent bracket structure of the BCH formula. Nevertheless, this experiment demonstrates the benefits of the proposed bracket non-linear layer.
@@
Vector Neurons 2.388 2.208 2.388 2.208
@@ 5.1.2 Learning Dynamics in Body Frame
$I\dot{\omega}(t)-\omega(t)\times I\omega(t)=M(t),$ (23)
where $\omega\in\mathbb{R}^{3}$ is the angular velocity, $I$ is the inertia tensor, and $M$ is the wrench input. It is an ordinary differential equation (ODE). Given the ODE and an initial condition $\omega_{0}$ at time $t_{0}$, we can solve the initial value problem, i.e., predict the trajectory of the system. In this experiment, we learn the ODE from historic trajectory data and predict the trajectories of the learned system given arbitrary initial conditions.
As a case study, we aim to learn the dynamics of the free-rotating (i.e., $M(t)=0$) International Space Station (ISS) from the National Aeronautics and Space Administration (2002).
Specifically, we can rewrite the ODE as the vector field
$\dot{\omega}=f(\omega;I)$ (24)
to represent the corresponding system dynamics, and we fit $f$ using a neural network. When a change of reference frame is performed, the inertia tensor undergoes a conjugation action. As a result, $f$ is equivariant under the change of reference frame: $\forall R\in\mathrm{SO}(3)$,
@@
We use the Neural ODE (Chen et al., 2018) framework to train the ODE from trajectory data. As shown in Figure 2, it consists of a neural network that models $f$ and an ODE solver. We compare our models with the model used in the original Neural ODE paper, which is an MLP network.
Figure 2: The framework used in modeling the Euler-Poincaré equation. The Lie neurons learn the dynamic equation, and an off-the-shelf ODE solver is employed to solve the ODE. Here, $m$ is a set of learnable equivariant weights.
@@
Using the inertia tensor from the National Aeronautics and Space Administration (2002), we randomly generate $10$ trajectories, each containing $25$ seconds of data and $1000$ data points. We then evaluate the trained model using $10$ unseen trajectories in the test set. To analyze the equivariance property, we rotate the test inputs and trajectories using $10$ random rotations, which we denote as $\mathrm{SO(3)}$ in Table 2. The table reports the norm distance between the estimated and the ground truth trajectories, evaluated at different points of time. We observe that the baseline is unable to predict the trajectories on the test set correctly. Therefore, we additionally report the results when both models are trained and evaluated on a single trajectory (The training and test sets are identical).
The proposed method obtains accurate predictions for all experiments. The prediction error remains low when the reference frames are changed. In comparison, the MLP is able to overfit on a single trajectory but fails to generalize when the reference frame is rotated. The MLP is also unable to correctly predict the trajectories when evaluated on the unseen data.
When the mixing layers is not added, the network is unable to converge. It is likely because in (23), the inertia tensor acts on the left of the angular velocity, which results in mixing in the geometric dimension. Without the mixing layers, such an operation cannot be modeled correctly, which demonstrates the value of the proposed mixing layers.
@@ 5.1.3 Point Cloud Registration
To better connect with previous literature, we implement the Lie Neurons in the point cloud registration task, which follows the setup in (Zhu et al., 2022). The network is trained and evaluated on ModelNet40, which contains 3D models of objects in 40 categories. We corrupt both the training and testing data using Gaussian noise with a standard deviation of $0.01$ after normalizing the point cloud to a unit cube.
We compare the Lie Neurons with the mixing module against the Vector Neuron (Deng et al., 2021). In addition, to demonstrate the benefits of geometric mixing, we integrate the mixing module in the original Vector Neuron to serve as an additional baseline.
Table 3 shows the average registration error in degrees. On average, the mixing module improves the performance of the Vector Neurons. The Lie Neurons perform similarly to the Vector Neurons with the mixing module. This experiment once again demonstrates the benefits of geometric mixing.
@@ 5.2 Experiments on $\mathfrak{sl}(3)$
We perform three experiments here. The first two are regressions of invariant and equivariant functions that take $\mathfrak{sl}(3)$ elements as input. Due to the page limit, we elaborate on them in the Appendix. The third experiment is a classification of Platonic solids based on the homography relationship among their faces. Across all three experiments, we compare our method with a standard 3-layer MLP by flattening the input to $\mathbb{R}^{K\times C\times N}$. In addition, we set the feature dimension to 256 for all models. The architecture of each model can be found in Figure 5 in the Appendix.
@@ 5.2.1 Platonic Solid Classification
Each network is trained in 5 separate instances to analyze the consistency of the method. For the detailed architecture, we again refer the readers to Figure 5. Table 4 shows the classification accuracy. The LN achieves higher accuracy than the MLP. Since the MLP is not invariant to the adjoint action, its accuracy drops drastically when the camera is rotated. We also notice that the LN-LB performs slightly worse than the other two formulations.
@@ 6 Discussion and Limitations
Lie Neurons is a group adjoint equivariant network by construction. It does not require the Lie group to be compact. However, the LN-ReLU layer relies on a non-degenerated Killing form. As a result, the current formulation can only operate on semisimple Lie algebras. Secondly, the Lie Neurons take elements on the Lie algebra as inputs, but most modern sensors return measurements in standard vector spaces. More practical applications of the proposed work are yet to be explored. Lastly, this work assumes a basis can be found for the target Lie algebra, which is valid for many robotics and computer vision applications.
@@ 7 Conclusion
We demonstrate the effectiveness of the proposed method on several applications, including regression of the BCH formula on $\mathfrak{so}(3)$, dynamic modeling of the free-rotating ISS, point cloud registration, and classification tasks on $\mathfrak{sl}(3)$. These experiments clearly show the advantages of an adjoint-equivariant Lie algebraic network. We believe the Lie Neurons could open new possibilities in both equivariant modeling and more general deep learning on Lie algebras.
Impact Statements
@@
Appendix A An Illustration of the Proposed Work
@@
Appendix B Network Architectures
Figure 5 shows the different architecture used in each experiment. For the point cloud registration tasks, we follow the structure in (Zhu et al., 2023a) and only add mixing layers before each ReLU.
@@ Appendix B Network Architectures
Appendix C Extra experiments on $\mathfrak{sl}(3)$
We present two additional experiments on $\mathfrak{sl}(3)$, in which we ask the network to regress an invariant and an equivariant function.
C.1 Invariant Function Regression
@@
Table 5: The mean squared errors and the invariant errors on the invariant function regression task. $\downarrow$ means the lower the better.
@@
[…]
C.2 Equivariant Function Regression

### v3

@@ 5.2 Experiments on $\mathfrak{so}(3)$
As discussed in Section 4.6, our network when applied on $\mathfrak{so}(3)$ resembles Vector Neurons (Deng et al., 2021), but contains more flexible component layers. In this section, we present our results on the Baker–Campbell–Hausdorff (BCH) formula regression, dynamics learning, and point cloud registration.
@@ 5.2.1 Baker–Campbell–Hausdorff Formula
The BCH formula provides a way to compute the product of two exponentials of elements in a Lie algebra locally (Hall, 2013). For $X,Y,Z\in\mathfrak{g}$, and
@@
We note that the BCH formula is adjoint equivariant. This formula is widely used in robotics and control (Yoon et al., 2023; Chauchat et al., 2018; Kobilarov & Marsden, 2011). However, the higher-order terms are often discarded in most applications, which can result in a significant drop in accuracy. An accurate modeling of the BCH formula can have great benefits in the above field.
We generate $10,000$ training and $10,000$ testing data points for both $X$ and $Y$ and use (20) to generate the ground truth value. To ensure an injective exponential function, we limit the angle to $[0,\pi)$. We design the loss function to be:
@@
where $\|\cdot\|_{\mathrm{F}}$ denotes the Frobenius norm, and $I$ is the identity matrix. The network architecture used in this experiment can be found in Figure 5 in the Appendix. In addition to the Frobenius norm, we also report the log error, which is defined as: $E_{log}=||\log(e^{X}e^{Y}e^{-f(X,Y)})^{\vee}||,$ where |||||$\cdot$||| | ⋅ | | is the standard vector norm on $\mathbb{R}^{3}$.
The results are presented in Table 2. We compare Lie Neurons with Vector Neurons (Deng et al., 2021), a steerable equivariant network e3nn (Geiger & Smidt, 2022), the Equivariant MLP (Finzi et al., 2021), and an MLP with augmentation. In addition, we report the results obtained using (21) by truncating to only first-third order terms. Both EMLP and our method achieve low estimation errors, while the error of our method is one order of magnitude smaller than EMLP. Although the MLP is able to produce acceptable results on the non-conjugated test set, its performance suffers from a significant drop when the inputs are adjoint-transformed. With training augmentation, the performance of MLP is improved on the testing augmentation but is slightly reduced on the non-conjugated case. The e3nn and Vector Neurons are unable to converge in this experiment. Since Lie Neurons specialize to Vector Neurons on $\mathfrak{so}(3)$ if only the ReLU layer is used, this experiment demonstrates the benefits of the proposed bracket non-linear layer.
@@
First Order Approx 0.629 0.464 0.629 0.464
Second Order Approx 0.338 0.247 0.338 0.247
Third Order Approx 0.191 0.136 0.191 0.136
@@
MLP Augmentation 0.025 0.018 0.280 0.221
EMLP (Finzi et al., 2021) $2.6\times 10^{-3}$ $1.8\times 10^{-3}$ $2.6\times 10^{-3}$ $1.8\times 10^{-3}$
e3nn (Geiger & Smidt, 2022) 0.641 0.471 0.641 0.471
Vector Neurons (Deng et al., 2021) 0.617 0.454 0.617 0.454
@@ 5.2.2 Learning Dynamics in Body Frame
$I\dot{\omega}(t)+\omega(t)\times I\omega(t)=M(t),$ (23)
where $\omega\in\mathbb{R}^{3}$ is the angular velocity, $I$ is the inertia tensor, and $M$ is the torque input. This equation is an ordinary differential equation (ODE). Given the ODE and an initial condition $\omega_{0}$ at time $t_{0}$, we can solve the initial value problem, i.e., predict the trajectory of the system. In this experiment, we learn the ODE from historic trajectory data and predict the trajectories of the learned system given arbitrary initial conditions.
As a case study, we aim to learn the dynamics of the free-rotating (i.e., $M(t)=0$) International Space Station (ISS) from the National Aeronautics and Space Administration (2002). Specifically, we can rewrite the ODE as the vector field $\dot{\omega}=f(\omega;I)$ to represent the corresponding system dynamics, and we fit $f$ using a neural network. When a change of reference frame is performed, the inertia tensor undergoes a conjugation action. As a result, $f$ is equivariant under the change of reference frame: $\forall R\in\mathrm{SO}(3)$,
@@
We use the Neural ODE (Chen et al., 2018) framework to train the ODE from trajectory data. As shown in Figure 2, it consists of a neural network that models $f$ and an ODE solver. We compare our models with the model used in the original Neural ODE paper, which is an MLP network. In addition, we replace the MLP network with EMLP (Finzi et al., 2021) to serve as a equivariant baseline.
Figure 2: The framework used in modeling the Euler-Poincaré equation. Lie neurons learn the dynamic equation, and an off-the-shelf ODE solver is employed to solve the ODE. Here, $m$ is a set of learnable equivariant weights.
@@
Using the inertia tensor from the National Aeronautics and Space Administration (2002), we randomly generate $10$ trajectories, each containing $25$ seconds of data and $1000$ data points. We then evaluate the trained model using $10$ unseen trajectories in the test set. To analyze the equivariance property, we rotate the test inputs and trajectories using $10$ random rotations, which we denote as $\mathrm{SO(3)}$ in Table 3. The table reports the norm distance between the estimated and the ground truth trajectories, evaluated at different points in time. We observe that the baselines are unable to predict the trajectories on the test set correctly. Therefore, we additionally report the results when both models are trained and evaluated on a single trajectory (The training and test sets are identical).
The proposed method obtains accurate predictions for all experiments. The prediction error remains low when the reference frames are changed. In comparison, both the MLP and EMLP are unable to correctly predict the trajectories when evaluated on the unseen data. When trained and tested on the same trajectory, both the MLP and EMLP can overfit the data. However, the MLP fails to generalize when the reference frame is rotated, while the EMLP slightly alleviates such a problem.
We observe that when the mixing layers are not added, the network is unable to converge. It is likely because, in (23), the inertia tensor acts on the left of the angular velocity, which results in mixing in the geometric dimension. Without the mixing layers, such an operation might not be modeled correctly, which demonstrates the value of the proposed mixing layers.
@@ 5.2.3 Point Cloud Registration
In this experiment, we follow the setup in (Zhu et al., 2022) and implement Lie Neurons in the point cloud registration task. The inputs are two noisy point clouds with different orientations. The goal of the network is to regress an $SO(3)$ rotation that best aligns the two point clouds in a correspondence-free manner. The network is trained and evaluated on ModelNet40, which contains 3D models of objects in 40 categories.
We compare Lie Neurons with the mixing module against Vector Neuron. In addition, to demonstrate the benefits of geometric mixing, we integrate the mixing module in the original Vector Neuron to serve as an additional baseline.
Table 4 shows the average registration error in degrees. On average, the mixing module improves the performance of Vector Neurons. Lie Neurons perform similarly to Vector Neurons with the mixing module. This experiment once again demonstrates the benefits of the geometric mixing layer.
@@ 5.3 Experiments on $\mathfrak{sl}(3)$
We perform three experiments for $\mathfrak{sl}(3)$: a classification of Platonic solids and 2 function regression tasks. We report the function regression tasks in Appendix D.
@@ 5.3.1 Platonic Solid Classification
Each network is trained in 5 separate instances to analyze the consistency of the method. For the detailed architecture, we again refer the readers to Figure 5. Table 5 shows the classification accuracy. The LN achieves higher accuracy than the MLP. Since the MLP is not invariant to the adjoint action, its accuracy drops drastically when the camera is rotated. When trained with augmented data, the MLP performance on the rotated test set increases, but the overall performance decreases. We also notice that the LN-LB performs slightly worse than the other two formulations.
@@ 6 Discussion and Limitations
Lie Neurons are a group adjoint equivariant network by construction. It does not require the Lie group to be compact. However, the LN-ReLU layer relies on a non-degenerated Killing form, which limits the operation on semisimple Lie algebras for such a layer. For general Lie groups, the adjoint representation might not be irreducible. As a result, the linear layer may not cover all equivariant maps. However, given the complex representation theory of semisimple groups, we use the current linear layer design and increase flexibility with various nonlinear layers. The current mixing module only works for $\mathfrak{so}(n)$. Nevertheless, we conjecture it is possible to extend to any semi-simple Lie algebra, as discussed in Section 4.3. Lie Neurons take elements in the Lie algebra as inputs, but most modern sensors return measurements in standard vector spaces. Finding equivariant lifts from the measurement space to the Lie algebraic space is an important future work. Lastly, this work assumes a basis can be found for the target Lie algebra, which is valid for many robotics and computer vision applications.
@@ 7 Conclusion
We demonstrate the effectiveness of the proposed method on several applications, including function regression on $\mathfrak{sp}(4)$, regression of the BCH formula on $\mathfrak{so}(3)$, dynamic modeling of the free-rotating ISS, point cloud registration, and classification tasks on $\mathfrak{sl}(3)$. These experiments clearly show the advantages of an adjoint-equivariant Lie algebraic network. We believe Lie Neurons could open new possibilities in both equivariant modeling and more general deep learning on Lie algebras.
Impact Statement
@@
Acknowledgments
This work was supported by AFOSR MURI FA9550-23-1-0400.
Appendix A Additional Preliminaries
A.1 The Hat and Vee Operator
As introduced in Section 3, the Hat and Vee operators can be defined as (Chirikjian, 2011):
[…]
where $E_{i}=\in\mathfrak{g}$ are linear independent basis in $\mathfrak{g}$, and $e_{i}$ are the canonical basis of $\mathbb{R}^{m}$.
With $\begin{bmatrix}0&-\omega_{z}&\omega_{y}\\ \omega_{z}&0&-\omega_{x}\\ -\omega_{y}&\omega_{x}&0\end{bmatrix}\in\mathfrak{so}(3)$, we have
$\small W$ $=\begin{bmatrix}0&-\omega_{z}&\omega_{y}\\ \omega_{z}&0&-\omega_{x}\\ -\omega_{y}&\omega_{x}&0\end{bmatrix}$ (26)
$=\omega_{x}E_{x}+\omega_{y}E_{y}+\omega_{z}E_{z}$ (27)
$=v^{\wedge},$ (28)
$W^{\vee}$ $=v=\begin{bmatrix}\omega_{x},\omega_{y},\omega_{z}\end{bmatrix}^{\mathsf{T}}.$ (29)
Appendix B An Illustration of the Proposed Work
@@
Appendix C Implementation Details
Figure 5 shows the different architecture used in each experiment. For the hyperparameter used in each experiment, we kindly refer the readers to the open-sourced GitHub repository: https://github.com/UMich-CURLY/LieNeurons.
C.1 Baker–Campbell–Hausdorff Formula
The Lie Neurons structure used in the BCH experiments is shown in Figre 5. It consists of two LN-LB+LN-LR layers. The feature dimension of each layer is set to $1024$, while the last linear layer projects the features back to dimension $3$.
For Vector Neurons, we maintain the same architecture as Lie Neurons while changing all the LN-LB to the ReLU layer from VN. The feature dimension is also set to $1024$.
For EMLP, we construct 3 EMLP blocks, each consisting of a linear layer, a bilinear layer, and a gated nonlinearity. The channel size is set to 128 for each block. (We were unable to increase the feature dimension due to memory complexity.) The input representation is set to two $3\times 1$ vectors with respect to $SO(3)$, and the output representation is set to one $3\times 1$ vector with respect to $SO(3)$.
To construct an e3nn network for comparison, we use equivariant linear layers, batch normalization layers, and activation layers from the library. We experiment with both norm-based activation layers and spherical point-wise activation layers. We use type-0 to type-4 features in the hidden layers. We tune the hyperparameters with different choices of depth and feature dimensions. The reported result is using 6 linear-batchnorm-norm-based-activation layers. The network size is comparable to other networks in the comparison. In our experiment, the performance of e3nn is similar to Vector Neurons, both of which cannot successfully regress the BCH formula. It is possible that e3nn might be able to achieve better performance by designing more complicated architectures, nonlinearities, and further tuning the hyperparameters, but the experiment still confirms the superiority of our proposed LN model in this equivariant regression task on so(3) algebra.
C.2 Dynamic Modeling
When a change of reference frame is performed to a dynamical system, $\omega$ undergoes a left rotation action, while the inertia tensor undergoes a change of basis action. That is $RIR^{-1}$. The equivariance of the system can be shown using the following derivation:
$f(\omega,I)$ $=-I^{-1}(\omega\times I\omega)=-I^{-1}[\omega]_{\times}I\omega,$ (30)
where $[\cdot]_{\times}$ means skew symmetric form. Here we are using the fact that the cross product can be replaced with a skew symmetric matrix multiplication. By using $[R\omega]_{\times}=R[\omega]_{\times}R^{-1}$, we can have
$f(R\omega;RIR^{-1})$ $=-RI^{-1}R^{-1}(R\omega)\times RIR^{-1}R\omega$ (31)
$=-RI^{-1}R^{-1}R[\omega]_{\times}R^{-1}RIR^{-1}R\omega$ (32)
$=-RI^{-1}[\omega]_{\times}I\omega=Rf(\omega,I)$ (33)
This means that the symmetry of the system itself is with respect to both $\omega$ and $I$. If we force the network to learn the dynamics using only $\omega$ as the input, we cannot correctly capture the symmetry of the system. As a result, we introduce $m$ as steerable and learnable weights as additional input elements. During test time, the weights $m$ are fixed. When we rotate the system, we apply the same rotation on $m$. Since $m$ is introduced to address equivariance with respect to the inertia tensor, it can be viewed as an implicit representation of the inertia tensor. As shown in Figure 5, the learnable weights $m$ pass through two branches of LN-LR layers. After such, they are mixed with the input data x in a bracket layer. For each layer, we set the feature dimension as $20$.
For the baseline EMLP (Finzi et al., 2021), we introduce the learnable weights $m$ as three $3\times 1$ vectors to the system input to serve as the implicit representation of the inertia tensor. We construct 3 EMLP blocks, each consisting of a linear layer, a bilinear layer, and a gated nonlinearity. The channel size is set to 128 for each block. The input representation is set to four $3\times 1$ vectors with respect to $SO(3)$ (a measurement $\mathbf{x}$ and three $3\times 1$ vectors for $m$.), and the output representation is set to one $3\times 1$ vector with respect to $SO(3)$.
C.3 Point Cloud Registration
We modified the open-sourced implementation provided in Zhu et al. (2022) for the point cloud registration task.  [^3] We modified the open-sourced code from https://github.com/minghanz/EquivReg. The inputs are two noisy point clouds with different orientations. The goal of the network is to regress an $SO(3)$ rotation that best aligns the two point clouds in a correspondence-free manner. We corrupt both the training and testing data using Gaussian noise with a standard deviation of $0.01$ after normalizing the point cloud to a unit cube. The initial rotation is sampled from (0,180) degrees randomly. Similar to Zhu et al. (2022), we use PointNet as the encoder backbone for the network, and we replace the Vector Neurons modules with the Lie Neurons modules. In addition to Lie Neurons, we test the effect of the mixing layers with Vector Neurons. For that, we add the mixing layers before each ReLU in the original Vector Neurons while keeping the rest of the structure untouched.
C.4 Platonic Classification
The architectures used in this experiment are shown in Figure 5. Since classification is an invariant task, an invariant layer is attached to the end of each model. The feature dimension is set to 256 for each layer of the models. We compare our method with a standard 3-layer MLP, in which we also set the feature dimension to 256.
@@ C.4 Platonic Classification
Appendix D Additional Experiments on $\mathfrak{sl}(3)$
We present two additional experiments on $\mathfrak{sl}(3)$, in which we ask the network to regress an invariant and an equivariant function. Across both experiments, we compare our method with a standard 3-layer MLP by flattening the input to $\mathbb{R}^{K\times C\times N}$. In addition, we set the feature dimension to 256 for all models.
D.1 Invariant Function Regression
@@
Table 6: The mean squared errors and the invariant errors on the $\mathfrak{sl}(3)$ invariant function regression task. $\downarrow$ means the lower the better.
@@
LN-LR $Id$ 66,562 1.30 $\times$ 10-3 3.24 $\times$ 10-5 1.30 $\times$ 10-3 3.25 $\times$10-5 3.60 $\times$ 10-4 5.48 $\times$ 10-5
LN-LB $Id$ 132,098 0.557 1.87 $\times$ 10-4 0.557 1.87 $\times$ 10-4 $\mathbf{1.43\times 10^{-5}}$ 1.42$\times$10-6
LN-LR + LN-LB $Id$ 263,170 $\mathbf{8.84\times 10^{-4}}$ 2.52$\times$10-5 $\mathbf{8.84\times 10^{-4}}$ 2.49$\times$10-5 4.00$\times$10-4 0
D.2 Equivariant Function Regression